I'm stuck on this question and i can't seem to solve it. can you please help
(y+6)/5-(y-6)/49=1 I need to solve for y but i am having major trouble with this

Answer by nerdybill(7384)

(y+6)/5-(y-6)/49=1
multiply both sides by 245 to get:
(y+6)49-(y-6)5=245
(49y+294)-(5y-30)=245
49y+294-5y+30=245
44y+324=245
44y=-79
y = -79/44